### v2

Abstract
Advances in AI portend a new era of sophisticated disinformation operations. While individual AI systems already create convincing—and at times misleading—information, an imminent development is the emergence of malicious AI swarms. These systems can coordinate covertly, infiltrate communities, evade traditional detectors, and run continuous A/B tests, with round-the-clock persistence. The result can include fabricated grassroots consensus, fragmented shared reality, mass harassment, voter micro-suppression or mobilization, contamination of AI training data, and erosion of institutional trust. With increasing vulnerabilities in democratic processes worldwide, we urge a three-pronged response: (1) platform-side defenses—always-on swarm-detection dashboards, pre-election high-fidelity swarm-simulation stress-tests, transparency audits, and optional client-side “AI shields” for users; (2) model-side safeguards—standardized persuasion-risk tests, provenance-authenticating passkeys, and watermarking; and (3) system-level oversight—a UN-backed AI Influence Observatory.
Keywords AI Swarms  $\cdot$ Artificial Intelligence  $\cdot$ Democracy  $\cdot$ Disinformation  $\cdot$ Influence Operations  $\cdot$ Misinformation
The Rise of Automated Influence
Public opinion manipulation has entered a dangerous new phase, amplifying and complicating its traditional foundations in rhetoric and propaganda. Rapid advances in large language models (LLMs) and autonomous agents [1, 2] now let influence campaigns reach unprecedented scale and precision. Leading AI researchers caution that AI could foster or magnify mass manipulation [3], and cross-national studies confirm that generative tools dramatically expand propaganda output without sacrificing credibility [4, 5, 6]. Controlled experiments likewise show most LLMs can inexpensively create election falsehoods that appear more human than real human text [5, 7, 8, 6]. Techniques meant to refine AI reasoning, such as chain-of-thought prompting [9, 10], can just as effectively polish convincing lies. In short, automated, scalable influence operations are reshaping the information landscape and pose significant risks for society.
Enabled by these capabilities, another disruptive shift is emerging: swarms of collaborative, malicious AI agents. Fusing LLM reasoning with multi-agent architectures [1], these systems are capable of coordinating autonomously, infiltrating communities, and fabricating consensus at minimal cost. Where legacy botnets acted like megaphones, repeating one script, AI swarms behave like adaptive conversationalists with thousands of distinct personas that learn from feedback, pivot narratives, and blend seamlessly into real discourse. By mirroring human social dynamics using adaptive tactics, they threaten democratic discourse.
Urgent action is needed before such systems become fully entrenched in our global information ecosystem, threatening the integrity of democratic processes worldwide. This policy brief situates the emerging AI swarm era within the broader evolution of digital influence operations, diagnoses its unprecedented risks, and sets out a policy agenda.
Evolution of Information Operations
Information operations long predate social media: Cold-War propaganda ranged from Soviet claims that AIDS was a US bioweapon to documented covert U.S. efforts against Latin-American leaders. Online manipulation then accelerated in the 2000s–2010s, targeting Brexit, the 2016 US election, and campaigns in Brazil and the Philippines, and even fueling violence such as the Rohingya genocide [11, 12, 13, 14, 15, 16]. Field experiments still find modest, if any, attitude shifts [17, 18], yet sustained state and private investment, along with documented links between hate campaigns and offline harm [19], ensure tactics will keep evolving.
The Russian Internet Research Agency’s 2016 Twitter operation shows the limits of manual botnets: one percent of users saw 70% of its content, with no measurable effects on opinions or turnout [20, 17, 18]. Human labor, blunt messaging, and slow iteration capped its reach. But large LLMs remove those caps. Readily available and frequently jailbreak-prone [21, 22], LLMs can now generate persuasive, tailored text at scale and have shifted deeply held beliefs in laboratory settings [23, 24]. Open-source releases by firms such as Meta and DeepSeek further lower access barriers.
Consequently, AI-supported election interference is no longer hypothetical. Taiwan’s and India’s 2024 campaigns saw AI-generated deepfakes, and entire fabricated news outlets now steer debates on topics from climate policy to foreign aid [25]. Absent guardrails, LLM-driven swarms could transform sporadic mis- and disinformation into persistent, adaptive manipulation of democratic discourse.
Swarm Capabilities
Recent breakthroughs in multi-agent systems (MAS) (i.e., computational environments where multiple AI agents interact, communicate, and collaborate to solve complex problems) have fused LLM reasoning with agentic memory, planning, and communication [26, 27, 28, 29]. Five technical advances now matter most for influence operations (see Figure 1).
A first capability is the shift from central command to a fluid swarm that coordinates in real time. A single adversary can operate thousands of AI personas, scheduling content and updating narrative frames instantaneously across an entire fleet of agents. Local adaptation plus periodic hub-sync blurs the line between command-and-control and emergent “hive” behavior. If these agent swarms evolve further into loosely governed “societies,” capable of internal norm formation and division of labor, the challenge shifts from tracing commands to understanding emergent group cognition [30, 31, 32, 29]. This evolution creates profound defensive challenges. Rather than simply tracing and blocking command sources, defenders must confront unpredictable collective behaviors, where new biases may emerge during social coordination that are not detectable at the individual agent level. More concerning, these societies may experience spontaneous or adversarially-induced norm shifts, abandoning their original engineered constraints for unpredictable new behavioral patterns through critical mass dynamics [33].
Second, agents now can map social graphs and slip into vulnerable communities with tailored appeals. Malicious agent swarm tactics include infiltrating vulnerable online communities to get humans to follow them and generating high volumes of content that appeals to those communities [34]. AI agents can employ systems that map social graphs at scale, identify key communities and beliefs, and track trending topics [35]. This can also work in a decentralized fashion while still having global, network–wide efficacy [36]. Equipped with such capabilities, agent swarms can position themselves for maximum impact and tailor messages to the beliefs and cultural cues of each community, enabling more precise targeting than previous botnet approaches.
Third, human-level mimicry lets swarms evade detectors that once caught copy-paste bots. Methods for detecting coordinated inauthentic behavior, such as influence operations, generally rely on activity patterns that are suspiciously similar between accounts and therefore statistically unlikely to result from independent behaviors [37, 38]. High-fidelity photorealistic avatars, context-aware slang, and heterogeneous posting rhythms circumvent the tell-tale synchrony that older detectors flag [39, 40].
Fourth, continuous A/B testing refines messages at machine speed. In the near term, human operators will still steer agent swarms while letting the agents perform limited in-context adaptation. However, over time, the balance is likely to tilt toward self-optimizing swarms that harvest their own reward signals: real-time engagement data, structural cues from recommender systems, or even natural-language critiques [41]. Even these systems cannot improve without such measurable feedback. Given sufficient signals, though, swarms can run millions of micro-A/B tests, propagate the winning variants at machine speed, and iterate far faster than human campaigners.
Finally, round-the-clock presence turns influence into a long-term, low-friction infrastructure. Unlike transient influence operations, agent swarms can persist for protracted durations, embedding themselves within communities over long timescales and gradually shifting discourse. This persistent presence amplifies the effectiveness of other mechanisms described above. In the realm of cognitive warfare, AI’s infinite, expendable machine time becomes a strategic weapon against the finite, emotionally taxing efforts of human engagement.
Figure 1: From Malicious AI Swarm to Democratic Harm. Concentric-circle schematic linking threat mechanics to civic outcomes. At the core, a Malicious AI Swarm supplies at least five key capabilities—decentralized orchestration, community infiltration, detection evasion, continuous optimization, and persistence (inner ring). These capabilities radiate outward into at least nine democratic-impact pathways, including synthetic consensus cascades, undermined collective intelligence, fragmented epistemic commons, LLM training-data poisoning, mass harassment, algorithmic overcompensation, FUD-induced disengagement, institutional legitimacy erosion, and engineered norm shifts and mass mobilization (outer ring). The upward arrow signals the causal flow from inner mechanics to systemic harm. Note that this flow is likely reciprocal, as impaired democratic functioning increases a society’s vulnerability to malicious AI swarms. FUD = Fear, Uncertainty, Doubt.
Pathways of Harms to Democracy
The unprecedented capabilities of AI swarm-driven influence campaigns pose profound threats to democracy by shaping public opinion formation [42, 43]. These capabilities convert into at least five cascading harms (see Figure 1).
The immediate effect is a synthetic consensus that exploits social-proof heuristics. Swarms may seed harmonized narratives across disparate niches, creating the illusion of majority agreement within days. Citizens update opinions based on perceived peer norms, not evidence. This chorus of seemingly independent voices creates a mirage of grassroots consensus with enhanced speed and persuasiveness. The result is a deeply embedded manipulation strategy that allows operators to nudge public discourse almost invisibly over extended periods of time.
This coordinated chorus also erodes the independence that collective intelligence requires, and which democracies rely on. While the previous effect exploits the influence of social norms, this second impact directly undermines human cognitive information processing systems. Specifically, the “wisdom of crowds” phenomenon, where aggregated judgments outperform individual experts, depends critically on independence between judgments [44, 45, 46]. When AI agents in a swarm coordinate to simulate diverse voices across platforms–appearing as distinct citizens, experts, and community leaders while actually representing a singular strategic objective–citizens can overestimate the informational value of this artificial consensus.
Collaborating AI agents can tailor streams of misleading information to each sub-community’s linguistic, cultural, and emotional markers, weaving partially overlapping but strategically segmented realities. Unlike passive “filter bubbles,” these engineered realities are designed to keep groups apart except when not desired, making cross-cleavage consensus increasingly unfeasible. Once initiated, such streams can also spread via social contagion, with the effect of bots potentially cascading beyond the people to whom the bots are directly connected, to two or three degrees of separation, as is often empirically the case [47].
By flooding the web with fabricated chatter, agent swarms can contaminate the data from which future models learn. Malicious operators spin up faux publics that flood the web. LLMs then ingest this chatter; at the next retraining cycle, fabricated narratives calcify inside the model weights [48]. Thus, AI swarms can rig the epistemic substrate on which future deliberation and future AI tools will rely, compounding democratic erosion.
Separate from fragmentation, swarms can at a low cost unleash coordinated synthetic harassment campaigns (i.e., “synthetic shitstorms”) that relentlessly target politicians, dissidents, whistleblowers, journalists, and their family and social networks with overwhelming, psychologically tailored abuse. Unlike conventional trolling, these operations appear as spontaneous public outrage while actually representing orchestrated action by thousands of AI personas adapting dynamically to target responses. By the time monitoring teams distinguish these AI campaigns from organic criticism, targets may have already withdrawn from public life, delivering significant victories for campaign operators while systematically excluding critical voices from democratic discourse.
Generally, as trust collapses, fear-uncertainty-doubt (FUD) can drive users into gated channels and silence. When citizens realize that vast portions of online speech may be AI-generated, baseline trust in social platforms and their users could sharply decline. Some threat actors may even welcome their synthetic interventions being exposed, reasoning that the mere revelation of manipulation can sow as much confusion as successful deception. Compounding this, genuine users will inevitably be misidentified as bots, weaponizing false accusations to discredit individuals and intensify FUD. This “epistemic vertigo” may mesh with a flood of low-cost LLM spam that overwhelms timelines, making genuine human conversation harder to find. Together, FUD and content saturation could drive users to disengage or retreat into gated channels, shrinking the shared public sphere on which deliberative democracy relies.
Algorithmic over-compensation can then elevate celebrity voices while sidelining ordinary citizens. When timelines flood with AI-authored posts, both ranking algorithms and users may retreat to obvious trust proxies—big follower numbers, blue checks, and legacy clout. Attention may concentrate around celebrities and major brands, while ordinary participants fade from view. The public sphere contracts from many-to-many dialogue to a few-to-many broadcast, a shift that erodes democratic pluralism and encourages further cynicism or migration to closed groups.
Swarms can even tip latent norms into action [49] or dampen conformity, accelerating anti-democratic political action [30]. Rather than occupying central or influential positions, these agents could operate on the periphery of social networks, where early mobilization often begins [50]. Similar strategies can be weaponized for micro-targeted voter suppression or mobilization on behalf of specific parties and causes. Reinforcement-learning agents run thousands of messaging experiments per hour, iteratively adjusting content while classifiers mine engagement metrics and textual responses to infer voting intent and gauge each tactic’s success.
Taken to extremes, coordinated doubt may corrode institutional legitimacy and invite “emergency” rule. By coordinating subtle yet increasingly profound doubts about electoral commissions, courts, or statistics bureaus, swarms could corrode procedural trust. As confidence falters, previously unthinkable “emergency” measures (e.g., postponing elections, rejecting certified results) may become palatable, especially when deep fake endorsements from fabricated civic leaders amplify the call [51, 5].
Governance Measures and Technical Defenses
The emergence of sophisticated, collaborative, and increasingly autonomous swarms of AI agents capable of orchestrating influence operations marks a critical juncture. The causal arrow runs both ways: while malicious swarms endanger democratic norms, the quality of democratic governance will in turn shape how potent, or containable, those swarms become. A full shift to continuous detection, simulation, and proactive, defense-oriented AI must therefore be implemented and operational before AI swarm capabilities reach mass availability.
The first line of defense is always-on detection of anomalous coordination, backed by public audits. To defend the public information sphere in real time, platforms and regulators require always-on detectors that scan live traffic for statistically anomalous coordination patterns—the digital fingerprints of inauthentic influence swarms. Here, advanced network analytics methods are needed that can (i) identify emergent clusters by surfacing camouflaged indicators of coordinated activity; and (ii) spot narrative-alignment drifts with streaming topic models and change-point analysis that flag lock-step shifts in sentiment [52]. Deployments require explicit democratic mandates, independent audits, and public transparency dashboards to guard against misuse [53].
To extend protection all the way to end users, platforms should expose the same risk signals through optional “AI shields”—lightweight browser or operating system modules that run on the client side. A shield labels posts that carry high swarm-likelihood scores, lets users down-rank or hide them, and surfaces short provenance explanations in situ. Because all scoring happens locally, the shield preserves privacy while giving citizens agency over their information diets. Aggregated, anonymized feedback from shields can also flow back into the public dashboard, creating a distributed early-warning mesh that strengthens the central detectors.
Complementing live monitors, high-fidelity simulation can stress-test those detectors before each election. Real-time monitors are effective only when they anticipate the tactics they will face. Because defenders rarely have access to the underlying code or even the decision-making logic of malicious swarms, high-fidelity, agent-based simulation may be the only reliable window into how these systems might behave. AI agents seeded into synthetic networks can replicate a platform’s graph structure, content cadence, and recommender logic, yielding traces that recalibrate live detectors. Iterative red-teaming then pits hostile swarms against these defenses, mapping persuasion ceilings and long-term persona strategies, and continuously hardening guardrails.
Where manipulation slips through, calibrated defensive agents can supply water-marked counter-narratives. A proactive defense strategy must harness AI technology itself to counteract manipulative influence operations. As threats evolve, defensive AI agents can strategically disseminate accurate information and pro-democratic counter-narratives, engage constructively with and warn targeted communities, and promote media literacy and resilience at scale  [23, 54]. That said, counter-messaging must opt for precision over volume: if defensive agents indiscriminately flood a platform, human voices could disappear in a sea of synthetic content, triggering the very collapse we seek to avert. Defensive AI should therefore intervene only where manipulation is detected, clearly watermark its output, and operate under human-in-the-loop oversight. Such calibrated use of AI can help build trust before crisis points emerge and ensure that credible, evidence-based information sources are embedded within public discourse, keeping online communities resilient without drowning out authentic participation.
Oversight must be democratic and global, anchored by a UN-backed AI Influence Observatory that combines early-warning detection methods with public certification of incidents. Building on proposals for technically proficient supervisory bodies [3] and the UN’s nascent AI Advisory Body [55], the Observatory could ingest anonymized coordination signals from participating platforms and model providers, fuse them with crowdsourced reports from civil-society labs, and issue rapid advisories ranking each incident by provenance confidence and projected civic harm. The Observatory should maintain and continually update an open, searchable database of verified influence-operation incidents, allowing researchers, journalists, and election authorities to track patterns and compare response effectiveness across countries in real time. To guarantee both legitimacy and skill, its governing board would mix rotating member-state delegates, independent technologists, data engineers, and civil society watchdogs.
Stronger provenance may help societies harden identity without muting speech as persuasive swarms proliferate [52]. Policymakers may incentivize rapid adoption of passkeys, cryptographic attestations, and federated reputation standards, backed by ongoing R&D to harden them against spoofing [56]. However, “proof-of-human” alone is no panacea: millions lack formal IDs, biometrics raise privacy risks, and verified accounts can be hijacked. Therefore, a key priority should be regulating the resulting commercial market that underpins large-scale manipulation, where private sellers offer services ranging from vanity metrics to coordinated influence operations at remarkably low costs.
Finally, frontier models must undergo standardized persuasion-risk evaluations and publish scores in model cards before release. Companies should disclose within 24 hours whether entities interacting with users are human or automated. Pre-bunking campaigns that educate citizens about AI tactics can build cognitive immunity, but paradoxically, the very alertness they foster can still be exploited by threat actors to deepen FUD. “Prosocial media” architectures on interoperable platforms can offer structural resilience [57]. Governments and tech firms must prioritize AI-safety R&D while funding independent assessments of misuse potential and societal impact.
In the coming years, democracies have a brief but crucial opportunity to pull AI-enabled influence operations back from the brink. If platforms deploy always-on swarm detectors, frontier labs submit models to standardized persuasion stress tests, and governments launch an AI Influence Observatory that publishes open incident telemetry, we can blunt the most destabilizing tactics before critical political future events, without freezing innovation. Doing so demands the same choreography already mastered by AI swarms: rapid iteration, transparent data-sharing, and tight coordination across scientific, civil-society, industry, and election-security teams. Like earlier calls to confront emerging bio-risks, success hinges on acting “collaboratively … without unnecessarily impeding scientific research,” while keeping the public square resilient and accountable [58]. By committing now to principled action, upcoming societal decisions could even become a proving ground—rather than a casualty—of democratic AI governance.
Acknowledgements
We would like to thank Jon Roozenbeek and Asbjørn Følstad for their valuable input to this paper. We are also grateful to Patrycja Antosz, Önder Gürcan, Ivan Puga Gonzalez, and Phil Tinn for the fruitful discussions that contributed to the development of this work.